OpenStudy (anonymous):

Solve: 2/3x+6 = x+2/x^2-10

OpenStudy (mertsj):

Cross multiply for starters.

OpenStudy (anonymous):

\[\frac{ 2 }{ 3x+6 } = \frac{ x+2 }{ x^{2}-10 }\]

OpenStudy (anonymous):

2x^2 - 20 = 3x^2 + 12 ?

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

Better use FOIL on (3x+6)(x+2)

OpenStudy (anonymous):

3x^2 + 2 + 6x + 12?

OpenStudy (mertsj):